# TUMLU: A Unified and Native Language Understanding Benchmark for Turkic Languages

### v1

@@ Abstract
Being able to thoroughly assess massive multi-task language understanding (MMLU) capabilities is essential for advancing the applicability of multilingual language models. However, preparing such benchmarks in high quality native language is often costly and therefore limits the representativeness of evaluation datasets. While recent efforts focused on building more inclusive MMLU benchmarks, these are conventionally built using machine translation from high-resource languages, which may introduce errors and fail to account for the linguistic and cultural intricacies of the target languages. In this paper, we address the lack of native language MMLU benchmark especially in the under-represented Turkic language family with distinct morphosyntactic and cultural characteristics. We propose two benchmarks for Turkic language MMLU: TUMLU is a comprehensive, multilingual, and natively developed language understanding benchmark specifically designed for Turkic languages. It consists of middle- and high-school level questions spanning 11 academic subjects in Azerbaijani, Crimean Tatar, Karakalpak, Kazakh, Tatar, Turkish, Uyghur, and Uzbek. We also present TUMLU-mini, a more concise, balanced, and manually verified subset of the dataset. Using this dataset, we systematically evaluate a diverse range of open and proprietary multilingual large language models (LLMs), including Claude, Gemini, GPT, and LLaMA, offering an in-depth analysis of their performance across different languages, subjects, and alphabets. To promote further research and development in multilingual language understanding, we release TUMLU-mini and all corresponding evaluation scripts [^1] https://github.com/ceferisbarov/TUMLU.
@@ 1 Introduction
In this paper, we focus on building a truly representative and inclusive single-language family benchmark to address previous problems and provide a challenging setting for LLM evaluation. TUMLU (Turkic Unified Multilingual Language Understanding) benchmark covers the following languages: Azerbaijani, Crimean Tatar, Turkish, Uyghur, Uzbek, Karakalpak, Kazakh, and Tatar. The dataset consists of 4-choice questions at middle- and high-school levels. It consists of 38139 questions across 8 languages and 11 subjects (see Figure 1 for a higher-level breakdown across languages). It is the first such benchmark to include Uyghur, Karakalpak, Tatar, or Crimean Tatar. It is also a significant improvement over existing benchmarks for Azerbaijani, Uzbek, and Kazakh. Turkish dataset is TurkishMMLU, which was a separate project Yüksel et al. (2024). The benchmark is also representative in terms of different scripts by including questions and answers in chosen languages in Latin, Cyrillic, and Arabic scripts. These datasets are transliterated such that it could be possible to get a dual dataset with the same content in two scripts for further comparative studies. We use these dual datasets to compare the performance of LLMs across different scripts.
@@ 2 Related Work
SeaEval was one of the first LLM benchmarks to include Turkish Wang et al. (2024a). Global MMLU contains Kyrgyz and Turkish subsets. INCLUDE contains Azerbaijani and Kazakh. MRL 2024 Shared Task on Multi-lingual Multi-task Information Retrieval Tinner et al. (2024) contains an Azerbaijani dataset, but it contains general language understanding tasks rather than world knowledge. Kardeş-NLU has introduced a multilingual language understanding benchmark Senel et al. (2024). But again, this benchmark contains general language understanding tasks that require no world knowledge. There are also monolingual benchmarks. Mukayese was one of the earliest general language understanding benchmarks in Turkish Safaya et al. (2022). TurkishMMLU and TR-MMLU Bayram et al. (2025) were the first native MMLU alternatives for the Turkish language. Another pilot study was performed to evaluate LLMs in Kazakh language Maxutov et al. (2024). While there are no peer-reviewed monolingual MMLU alternatives for Azerbaijani, there is a general language understanding benchmark Isbarov et al. (2024).
@@ 5 Results
Model Mean aze crh kaa kaz tat tur uig uzb
Claude 3.5 Sonnet 79.0 84.4 81.0 75.3 83.0 84.0 84.9 70.9 68.6
GPT-4o 75.1 82.4 70.5 70.8 81.0 80.4 83.7 66.5 65.4
Gemini 1.5 Pro 66.7 74.7 59.7 67.4 78.3 77.3 59.1 55.8 61.1
Gemini 1.5 Flash 64.6 72.1 67.5 61.2 68.4 66.7 73.1 57.4 50.0
Claude 3.5 Haiku 63.3 70.6 61.7 54.9 69.9 67.3 77.6 54.4 50.0
Qwen2.5 72B 61.5 70.1 61.8 54.6 62.4 62.5 73.9 56.0 50.4
Llama 3.3 70B 57.6 66.0 58.7 49.2 59.7 69.5 68.1 45.6 44.1
Gemma 2 27B 51.5 58.1 49.8 47.6 58.4 54.7 64.3 42.2 36.9
Llama 3.1 70B 50.9 67.6 56.1 49.0 47.7 54.6 64.7 28.5 39.0
Gemma 2 9B 46.8 53.7 46.8 40.8 49.1 51.8 60.1 35.8 36.1
Qwen2.5 7B 42.1 48.0 42.6 37.2 45.0 40.5 55.6 33.4 34.6
Llama 3.1 8B 39.7 48.4 35.7 33.4 46.4 44.1 44.4 35.0 29.9
Table 2: Average 5-shot performance of models on Azerbaijani (aze), Crimean Tatar (crh), Kara-Kalpak (kaa), Kazakh (kaz), Tatar (tat), Turkish (tur), Uyghur (uig), and Uzbek (uzb) datasets.
@@
Turkish 84.0 87.0 94.0 91.0 77.0 76.0 85.0
Uyghur 73.0 66.0 - - 75.8 66.0 73.5
Uzbek 69.0 73.0 64.0 68.0 70.0 55.0 81.0
@@
Model Mean aze kaz tur uzb
Claude 3.5 Sonnet 82.2 87.1 (+2.7) 84.1 (+1.1) 87.4 (+1.5) 70.1 (+1.5)
GPT-4o 78.4 82.9 (-0.5) 80.7 (-0.3) 84.0 (+0.3) 66.0 (+0.6)
Gemini 1.5 Pro 69.4 74.6 (-0.1) 74.7 (-3.6) 75.0 (+15.9) 53.4 (-7.7)
Claude 3.5 Haiku 67.0 77.0 (+6.4) 74.0 (+4.1) 70.7 (-6.9) 46.3 (-3.7)
Qwen2.5 72B 66.1 72.1 (+2.0) 63.9 (+1.5) 78.4 (+4.5) 49.9 (-0.5)
Gemini 1.5 Flash 65.1 73.6 (+1.5) 67.6 (-0.8) 69.4 (-3.7) 50.0 (0)
Llama 3.3 70B 59.5 69.9 (+3.9) 69.3 (+9.6) 75.9 (+7.8) 23.0 (-21.1)
Gemma 2 27B 58.0 62.9 (+4.8) 61.6 (+3.2) 66.4 (+2.1) 41.1 (+4.2)
Llama 3.1 70B 52.8 59.0 (-8.6) 61.7 (+14.0) 73.3 (+8.6) 17.3 (-21.7)
Gemma 2 9B 51.4 57.1 (+3.4) 52.6 (+3.5) 62.3 (+2.2) 33.7 (-2.4)
Qwen2.5 7B 47.1 48.0 (0.0) 46.4 (+1.4) 56.3 (+0.7) 37.9 (+2.3)
Llama 3.1 8B 36.4 40.7 (-7.7) 38.9 (-7.5) 45.0 (+0.6) 20.9 (-9.0)

### v2

@@ Abstract
Being able to thoroughly assess massive multi-task language understanding (MMLU) capabilities is essential for advancing the applicability of multilingual language models. However, preparing such benchmarks in high quality native language is often costly and therefore limits the representativeness of evaluation datasets. While recent efforts focused on building more inclusive MMLU benchmarks, these are conventionally built using machine translation from high-resource languages, which may introduce errors and fail to account for the linguistic and cultural intricacies of the target languages. In this paper, we address the lack of native language MMLU benchmark especially in the under-represented Turkic language family with distinct morphosyntactic and cultural characteristics. We propose two benchmarks for Turkic language MMLU: TUMLU is a comprehensive, multilingual, and natively developed language understanding benchmark specifically designed for Turkic languages. It consists of middle- and high-school level questions spanning 11 academic subjects in Azerbaijani, Crimean Tatar, Karakalpak, Kazakh, Kyrgyz, Tatar, Turkish, Uyghur, and Uzbek. We also present TUMLU-mini, a more concise, balanced, and manually verified subset of the dataset. Using this dataset, we systematically evaluate a diverse range of open and proprietary multilingual large language models (LLMs), including Claude, Gemini, GPT, and LLaMA, offering an in-depth analysis of their performance across different languages, subjects, and alphabets. To promote further research and development in multilingual language understanding, we release TUMLU-mini and all corresponding evaluation scripts [^1] https://github.com/ceferisbarov/TUMLU.
@@ 1 Introduction
In this paper, we focus on building a truly representative and inclusive single-language family benchmark to address previous problems and provide a challenging setting for LLM evaluation. TUMLU (Turkic Unified Multilingual Language Understanding) benchmark covers the following languages: Azerbaijani, Crimean Tatar, Turkish, Uyghur, Uzbek, Karakalpak, Kyrgyz, Kazakh, and Tatar. The dataset consists of 4-choice questions at middle- and high-school levels. It consists of 38139 questions across 8 languages and 11 subjects (see Figure 1 for a higher-level breakdown across languages). It is the first such benchmark to include Uyghur, Karakalpak, Tatar, or Crimean Tatar. It is also a significant improvement over existing benchmarks for Azerbaijani, Kazakh, Kyrgyz, and Uzbek. The Turkish dataset is TurkishMMLU, which was a separate project Yüksel et al. (2024). The benchmark is also representative in terms of different scripts by including questions and answers in chosen languages in Latin, Cyrillic, and Arabic scripts. These datasets are transliterated such that it could be possible to get a dual dataset with the same content in two scripts for further comparative studies. We use these dual datasets to compare the performance of LLMs across different scripts.
@@ 2 Related Work
SeaEval was one of the first LLM benchmarks to include Turkish Wang et al. (2024a). Global MMLU contains Kyrgyz and Turkish subsets. INCLUDE contains Azerbaijani and Kazakh. MRL 2024 Shared Task on Multi-lingual Multi-task Information Retrieval Tinner et al. (2024) contains an Azerbaijani dataset, but it contains general language understanding tasks rather than world knowledge. Kardeş-NLU has introduced a multilingual language understanding benchmark Senel et al. (2024). But again, this benchmark contains general language understanding tasks that require no world knowledge. There are also monolingual benchmarks. Mukayese was one of the earliest general language understanding benchmarks in Turkish Safaya et al. (2022). TurkishMMLU and TR-MMLU Bayram et al. (2025) were the first native MMLU alternatives for the Turkish language. Another pilot study was performed to evaluate LLMs in Kazakh language Maxutov et al. (2024). KazMMLU is another monolingual MMLU-style benchmark for the Kazakh language Togmanov et al. (2025). While there are no peer-reviewed monolingual MMLU alternatives for Azerbaijani, there is a general language understanding benchmark Isbarov et al. (2024).
@@ 5 Results
Model mean aze crh kaa kaz kir tat tur uig uzb
Claude 3.5 Sonnet 78.9 84.4 81.2 75.3 83.0 75.7 84.0 85.7 71.3 69.1
GPT-4o 74.9 82.4 70.5 70.8 81.0 72.9 80.5 83.7 66.5 65.4
Gemini 1.5 Pro 73.8 78.6 70.3 68.2 78.4 72.3 80.5 80.0 71.0 65.1
Gemini 1.5 Flash 65.4 72.4 68.0 61.2 68.6 63.2 68.3 76.6 57.8 52.1
Claude 3.5 Haiku 64.0 70.6 62.9 55.2 69.9 64.8 67.5 78.0 56.6 50.3
Llama- 3.1 405B 63.0 65.9 69.5 60.0 69.0 64.1 70.4 59.7 58.2 50.4
Qwen2.5 72B 59.9 70.1 61.8 54.6 62.6 47.5 62.5 73.9 56.0 50.4
Llama 3.3 70B 58.6 66.0 58.7 49.2 60.0 60.2 69.5 68.4 51.6 44.1
Llama 3.1 70B 57.7 68.1 57.3 49.9 56.4 58.4 66.2 64.9 52.4 45.3
Gemma 2 27b 51.8 58.1 49.8 47.6 58.4 53.4 54.9 64.3 42.2 37.6
Gemma 2 9b 47.1 53.7 46.8 40.8 49.1 48.9 51.8 60.4 35.8 36.1
Qwen2.5 7B 41.0 48.0 42.6 37.2 45.0 31.8 40.5 55.6 33.4 34.6
Llama 3.1 8B 40.3 48.4 35.7 33.4 46.4 41.8 44.1 47.7 35.0 29.9
Table 2: Average 5-shot performance of models on Azerbaijani (aze), Crimean Tatar (crh), Karakalpak (kaa), Kazakh (kaz), Kyrgyz (kir), Tatar (tat), Turkish (tur), Uyghur (uig), and Uzbek (uzb) datasets.
@@
Turkish 84.0 88.0 94.0 92.0 78.0 79.0 85.0
Uyghur 75.0 66.0 - - 75.8 66.0 73.5
Uzbek 71.0 73.0 64.0 70.0 70.0 55.0 81.0
@@
Model Mean aze kaz tat tur uzb
Claude 3.5 Sonnet 84.0 87.1 (+2.7) 84.1 (+1.1) 87.9 (+3.9) 87.9 (+2.1) 72.9 (+3.7)
GPT-4o 79.4 82.9 (+0.4) 80.7 (-0.3) 83.0 (+2.5) 84.0 (+0.3) 66.3 (+0.9)
Gemini 1.5 Pro 76.6 80.0 (+1.4) 75.1 (-3.3) 79.9 (-0.6) 81.0 (+1.0) 67.0 (+1.9)
Llama 3.1 405B 68.9 73.4 (+7.6) 68.7 (-0.3) 65.0 (-5.4) 80.7 (+21.0) 56.4 (+6.0)
Claude 3.5 Haiku 70.0 77.0 (+6.4) 74.0 (+4.1) 72.2 (+4.8) 77.6 (-0.4) 49.0 (-1.3)
Gemini 1.5 Flash 68.1 73.9 (+1.4) 69.0 (+0.4) 70.1 (+1.8) 73.6 (-3.0) 54.1 (+2.0)
Qwen2.5 72B 67.1 72.1 (+2.0) 63.9 (+1.3) 67.6 (+5.1) 78.4 (+4.6) 53.6 (+3.1)
Llama 3.3 70B 66.8 70.6 (+4.6) 69.3 (+9.3) 66.3 (-3.2) 77.4 (+9.0) 50.4 (+6.3)
Gemma 2 27B 59.4 63.0 (+4.9) 61.6 (+3.1) 61.0 (+6.1) 66.4 (+2.1) 44.9 (+7.3)
Llama 3.1 70B 56.2 59.4 (-8.7) 61.7 (+5.3) 58.6 (-7.6) 73.3 (+8.4) 27.9 (-17.4)
Gemma 2 9B 52.0 57.3 (+3.6) 52.7 (+3.6) 50.1 (-1.7) 62.3 (+1.9) 37.4 (+1.3)
Qwen2.5 7B 46.4 48.1 (+0.1) 46.4 (+1.4) 43.0 (+2.5) 56.3 (+0.7) 38.0 (+3.4)
Llama 3.1 8B 38.2 40.7 (-7.7) 38.9 (-7.6) 39.7 (-4.4) 45.1 (-2.6) 26.6 (-3.3)